January 14, 2008

Project Status: 14 January 2008 2030UTC

Long Wait Ahead

Weather Status: Wave approaching coast of NW US will brush the ARB with little chance of producing weather.Beyond this, the area looks to be ruled by dry conditions until 20-22 Jan timeframe when the arctic cold outbreak that was mentioned yesterday still shows a possibility of wringing out some moisture. Today the models are pretty stingy on bringing significant snow to the region. Along with this, the off shore ridge appears to adopt a strong positive tilt that directs any Pacific moisture rivers well up into AK and associated troughs move up the ridge and then down into central Canada keeping them well away from the ARB.

Summary: No IOP for the next 5 days.

Observation Summary: An RH sensor at Huysink appears to be bad and will probably be absent for the duration of the exercise. Bodega Bay's temperature and humidity sensor are out of action, but will be replaced with the next maintenance visit. Otherwise all sites appear to be up and running.

Decision Summary: No operations for the next 5 days.

Product Summary: Moisture products from Huysink and temperature and moisture products from Bodega Bay have been removed from the web page; for the latter, only temporarily. All HMT model products are operative. Workstations are likely to ship tomorrow to MRY and RNO.

McGinley, GSD

January 13, 2008

Project Status: 13 January 2008 2030UTC

Only Fog in the Valley; Long IOP Wait Ahead

Weather Status: Not much to update from yesterday. Foggy conditions were over most of the Central Valley. Next event of any sort is the tail end of a trough moving far to the the north on Jan 15. NWP not at all excited about this. With high amplitude ridge off the coast the first opportunity will be a 21 Jan trough coming from the north as the ridge retrogrades . Long range guidance is showing this digging off the coast and bringing some dynamics to the ARB area. However, conditions are likely to be cold with no sub-tropical connection.

Summary: NO IOP weather expected for the next 5 days.

Observational Status: All systems up and running.

Decision Summary: No operations for at least 5 days.

Product Summary: All products up and available.

McGinley, GSD

January 12, 2008

Project Status: 12 January 2008 2030UTC

A Few Light Showers This Morning in the ARB

Weather Status: A few light showers were active in the ARB area this morning with 0.14 inches at BLU between 12 Jan 14-18UTC (06-12L). This was a bit heavier than the models were predicting yesterday. The weak wave that was forcing these showers was producing more precipitation to the north. This should all be ending in the next few hours. The next minor wave brushes by to the north on Monday 15 Jan with only insignificant impacts for northern CA. Beyond that things look bleak with a closed off low out about 165W and a blocking ridge to the north. The ridge will retrograde slowly and become full-latitudinal, eventually allowing shortwaves to dive southward along the coast and at at least bring some dynamics to the ARB in the 20-22 Jan time frame.

Summary: No IOP for the next 5 days.

Observational Summary: All non-manned sites were up and running including BLU where power was restored.

Decision Summary: No operations for the next 5 days.

Product Summary: All products available.

McGinley, GSD


January 11, 2008

Project Status: 11 January 2008 2030UTC

All's Quiet in the West

Weather Status: Storm track continues to hold to the north. A minor wave will move through N. CA. on Saturday just brushing the ARB area. The NAM predicts no precipitation in the ARB while the GFS is bringing in a few hundredths. This will not qualify as an event. Most activity stays off the coast after that. 5-day forecasts drag polar air into the Great Basin under strong northerly flow with blocking ridge building to west. No activity likely for 10 days. Long range outlook from NWS visitor Ed Berry and Klaus Weickmann of PSD indicate a retrogression of the ridge is likely as the MJO moves into a position favoring troughs off the West Coast and a more active period after a 2 week hiatus in activity.

Summary: No IOP expected for at least 5 days.

Observational Summary: Good news...Cazadero is back on line as of 11 Jan 17UTC. BLU surface site is still out after being up on 9 Jan (went down at 9/19UTC).

Decision Summary: No operations for next 5 days.

Product Summary: Products appear to be up and running except as noted above. Models have been made more robust with automatic restarting procedures if one or more gets hung up.

McGinley, GSD

January 10, 2008

Project Status: 10 January 2008 2030UTC

Few Light Showers in ARB

Weather Status: A few light showers were in the central Sierras today with rain amounts from a few 100ths to 0.15 in. Most showers have moved generally south of the ARB and continue to drift SE. BLU Airport temperature is 36F so the event likely was rain. The next opportunity for precipitation in the ARB is Saturday into Sunday. Forecast consensus was for this being a weak event with precip (likely to be rain) in the GFS and ECMWF less than 0.1in. Long range outlook indicates a mostly dry central California through 10 days. A closed-off low that may develop off the coast is the only precipitation possibility that could change this outlook. Weickmann and Berry will provide an outlook for the next 2-4 weeks at tomorrow's telecon.

Summary: No IOP for the next 5 days.

Observational Summary: Crews are attempting to get power restored to Cazadero. For some reason BLU surface station went down again. Otherwise all automated sites appeared to be working.

Decision summary: No operations for the next 5 days.

Product Summary: Model products up and running. HMT products available for all active sites.

McGinley, GSD

January 9, 2008

Project Status: 9 January 2008 2030UTC

IOP#5 Comes to an End

Weather Status: The fast moving storm of yesterday barreled though the ARB from late morning 8 Jan 17UTC (09L) to early this morning 9 Jan 10UTC (02L). Precipitation amounts around the area were 1 to 2 inches of liquid, about 10 inches of snow at BLU. BLU remained in snow the entire period, getting a dose of ice at one point when supercooled cloud put some ice on the dish. Snow to liquid ratios were about 6 to1 becoming wetter in the latter part of the event (5 to 1). The freezing level that started out at about 3000ft seemed to rise through the event as evidenced by the increase from 3000 to 5000 ft at Chico. By the end of the event the temperature at BLU was -1 to 0C (30-31F)so it was likely the melting layer was just below. The next wave due to arrive tonight and Thursday continues to look weak on the numerical guidance. GFS and NAM estimates are 0.5 inches or so. There is an upper level jet over the ARB so increased amounts in bands might be expected. A weak trough moving north of the area on Saturday could bring scattered showers to the central Sierras, but nothing warranting operations. Beyond that GFS and ECMWF ensemble members are cutting off a low in the east Pacific that may drift east. This will bear watching next week. All in all pretty quiet for the next 10 days.

Summary: No system of interest for the next 5 days.

Observational Status: Sondes successfully released at SHS, RNO, and OAK from 8 Jan 12UTC (04L) to 9 Jan 03 UTC (19L) at 3-hour intervals; Radar was operational from 8 Jan 1530UTC (0730L) to 9 Jan 05UTC (21L). Radar crew departed the site at 8/22L and are returning to Boulder today. BLU surface site was down for the IOP but came up today; Lee Vining surface site down. Cazadero profiler down.

Decision Summary: No operations planned for the next 5 days.

Product Status: NWP high resolution guidance did well on QPF, but had the precip into the ARB a bit early. National NWP guidance was very helpful with good QPF and timing. ALPS Workstation news: GSD personnel finishing up work on software. Workstation will be shipped to MTY and RNO Friday (11Jan); ALPS software for SAC will be tarred and FTP'd Friday.

McGinley, GSD

Project Status: 9 January 2008 0100UTC (IOP#5 Update)

Report from BLU

Contacted radar crew. Had a number of heavy snow bands, last one was 8/23-9/00UTC. Sergey estimated 9 new inches with a 6 to 1 snow to water ratio. Thus 1.5 inches of liquid.....looks like good event. Radar dish is now covered with 3-4mm of ice due to an encounter with supercooled cloud. DAX radar indicates showers moving eastward with back edge over Central Valley so they should end in a couple of hours. Heaviest are now east and south; only scattered showers remain near BLU. Surface front looks to be in Central Valley so a 9/02-03UTC passage looks good. Snow plow operator will plow road from 19-21L so radar crew can leave the site tonight. Plan is to shut down radar at 9/04UTC (20L) lock things up and head to Auburn.

McGinley, GSD

January 8, 2008

Project Status: 8 January 2008 2030UTC

IOP#5 Underway

Weather Status: The rain/ snow bands associated with the IOP#5 upper level wave arrived a bit later than expected, moving into the BLU area at 8 Jan 17UTC (11L). BLU reported heavy snow at about 19UTC (11L) and this has been continuing. Maximum precipitation rates looked to be about 0.6 inches/3 hours. Freezing levels starting at 3500-4000ft should rise a bit but BLU should remain in snow for the duration. Coarse grid models were still supporting over an inch of liquid for the event with higher resolution model still going for 1.5 to 2 inches. With multiple precip bands lined up west of the ARB it looked pretty certain that the one-inch-plus criteria would be realized. Upper trough passage is scheduled for 9 Jan 02UTC (8/18L). Precipitation showers should hang on for a few hours after trough passage. End of the event looks to be 9 Jan 06UTC (8/22L). Next wave upstream continues to favor a more northerly trajectory, basically just brushing by the ARB. Models were generally forecasting a total of 0.3 - 0.5 inches of precip, so the event continues to look weak, and not justifying an IOP. Beyond this set of disturbances the outlook is for generally dry weather with only slight shower possibilities Saturday, and nothing much beyond that. Some ensemble solutions for next week indicate a closed low off the coast that could impact the ARB if it chooses to move east.

Summary: Precipitation and IOP#5 ending between 9 Jan 03-06UTC (8/19-22L). Second wave too weak for IOP.

Observational Status: SHS and special RAOBS were initiated at 8 Jan 12UTC (04L) at the 3 sites. Radar went into operation 8 Jan 1530UTC (0730L). Crew was reporting moderate snow at BLU at telecon time. Big Bend surface observation was back on line and reported 6 inches of snow and 0.3 inches of liquid in 1.5 hours (since 8 Jan 1730UTC [0930L]). Chico profiler was back up yesterday afternoon. Cazadero was still out, as was BLU's surface site. Radar crews may have to remain on site overnight since access road has 1 foot of snow on it and at least a foot more is possible.... plowing isn't likely until early Wed. Radar crews will secure site and return to ESRL on Wed afternoon.

Decision Summary: Last sonde release 9 Jan 03UTC (19L) all sites. Radar continues until precipitation terminates sometime between 9 Jan 03-06UTC (8/19-22L)

Product Status:One HMT ensemble member had to be started manually when the system failed for that run. More products were available on the HMT web site as power is being restored at instrument sites. BLU power may be out until this weekend. ALPS workstation systems being shipped to RNO and MRY and software being transferred to SAC, so we may have ALPS for the next IOP.

McGinley,GSD

January 7, 2008

Project Status: 7 January 2008 2030UTC

IOP#4 Over; IOP#5 Starts Tomorrow

Weather Status: Only light snow this morning at BLU with another 0.44 inches since yesterday. This brought the estimated total precipitation to 7 to 8 inches in the ARB...a great event. A new fast moving wave due to start impacting the ARB tomorrow will bring a bit more than an inch over a 12-18 hour period on Tuesday. This wave was still tapping a tropical IPW plume, but the speed of the wave was reducing the expected total precipitation. Rain looks like it will begin Tuesday early morning ( 8 Jan 12UTC), the the heaviest period from 8/21 to 9/02UTC (8/13L to 18L). Coarser grid models were predicting a little over an inch with higher resolution models between 1 and 2 inches. HMT models averaged about 1.5 inches. The wave certainly fits within the IOP criteria and we offically declare IOP#5 underway tomorrow (8 Jan 12UTC [04L]). A second wave moving in Wed night into Thurs, looked a lot weaker in the NWP products than yesterday. Most of the dynamics associated with storm look to be moving further north and hence precipitation predicted is below IOP criteria. We will look at this storms potential tomorrow to see if it has better IOP possibilities. Beyond this wave a ridge building over the west coast will keep activity to a minimum at least through the next 10 days.

Summary: IOP#5 starts 8 Jan 12UTC (04L); precipitation ending 9 Jan 06UTC (22L). Second wave will arrive late Wed/Thurs but is likely to be weak. No activity for an extended period beyond that.

Observational Status: Profiler sites at Chico and Cazadero are out due to power failure. Severe damage to powerlines will keep these sites out of action for another few days. BLU power is still out at the surface site but the observation at the airport was back on line today. Radar crew will deploy to radar and prepare site for IOP#5. A decision about establishing generator power to bring back the HMT instrumentation will be made later today after assessment by the radar crew. For IOP#5 sounding operations will commence Tuesday morning (8 Jan 12UTC [04L]) with the final sonde at 9 Jan 03UTC (8/19L). Radar will be operational from Tuesday morning through Tuesday night.

Decision summary: IOP#5 Radar and RAOB operations will begin at 8 Jan 12UTC (04L). Sondes will be every 3 hours ending with the release on 9 Jan 03UTC (8/19L). Radar operations will terminate 9 Jan 06UTC (8/22L).

Product Status: Other than the sites without power products were available; all HMT model products were available on the web.

McGinley, GSD

Project Status: 7 January 2008 0200UTC IOP#4 Update

The HYDROX radar site has been visited by the snow removal crew. Current snow depth on the access road is about 7-8 inches. This will likely be removed early Monday after additional snow tonight. Snow depth at the airport was about 2 feet. Radar crew can deploy as planned Monday. Crew will need 5-6 hours on site to get radar set up for IOP#5. Duane will set up generator to restore surface observations after coordinating with Clark King tomorrow. Report from Duane Hazen who drove up to Blue Canyon this afternoon.

McGinley, GSD